# A Unified View of Optimal Kernel Hypothesis Testing

### v1

@@ Abstract
This paper provides a unifying view of optimal kernel hypothesis testing across the MMD two-sample, HSIC independence, and KSD goodness-of-fit frameworks. Minimax optimal separation rates in the kernel and $L^{2}$ metrics are presented, with two adaptive kernel selection methods (kernel pooling and aggregation), and under various testing constraints: computational efficiency, differential privacy, and robustness to data corruption. Intuition behind the derivation of the power results is provided in a unified way accross the three frameworks, and open problems are highlighted.
@@
In Section 1, we formalise the hypothesis testing framework, highlighting test level and power properties. In Section 2, we present multiple testing via the aggregation method, which is strictly more powerful than using a Bonferroni correction. In Section 3, we consider hypothesis testing under three constraints: computational efficiency, differential privacy and robustness to data corruption. In Sections 4, 5 and 6, we introduce the two-sample, independence and goodness-of-fit testing frameworks, respectively, and derive power guarantees under kernel and $L^{2}$ uniform separation in the standard, efficient, private and robust settings, with aggregation and pooling kernel adaptation methods. In Section 7, we highlight open problems which are left for future work. In Appendix A, we provide intuition behind the proofs of all the power results presented.
@@ 1 Hypothesis testing
We formally introduce the concept of hypothesis testing in Section 1.1, we discuss test level control in Section 1.2 and test power guarantees in Section 1.3.
@@ 1.1 Definition of hypothesis testing
For two-sample testing, the space ${\mathcal{P}}$ consists of all pairs of distributions, and ${\mathcal{P}}_{0}$ of only the pairs with the same distribution for each component. For independence testing, the space ${\mathcal{P}}$ consists of joint distributions, and ${\mathcal{P}}_{0}$ of the ones which are equal to the product of their marginals. For goodness-of-fit testing, we have a reference model distribution $P_{\textrm{model}}$, the space ${\mathcal{P}}$ consists of all distributions while ${\mathcal{P}}_{0}=\{P_{\textrm{model}}\}$ is simply the model distribution. For more details on these testing frameworks, see Sections 4, 5 and 6.
@@
A test is a function which takes as input the i.i.d. samples from $P\in{\mathcal{P}}$, and returns 0 if the null ${\mathcal{H}}_{0}$ is believed to hold, or 1 otherwise (i.e. ${\mathcal{H}}_{1}$ is believed to hold). A test is usually constructed by first computing a statistic, a real value computed from the data which is designed to capture evidence against the null when it exists. This statistic is then compared to a rejection region to decide whether to reject the null. Equivalently, the test can be defined via its p-value rather than via its quantile (see details below in Section 1.2).
@@ 1.2 Level of hypothesis testing
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}(\textrm{reject }{\mathcal{H}}_{0})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \alpha.$ (2)
Typically, we would want this inequality to be tight: if the type I error is required to be less than $\alpha$, we would ideally want it to be as close as possible to $\alpha$ since a smaller type I error (e.g., conservative test) would result in a larger type II error (see Section 1.3).
@@
A sample is said to be exchangeable if, for any permutation, the joint distribution over the permuted samples is equal to the joint distribution over the original samples. Some hypothesis testing problems can be framed as testing whether exchangeability holds (Lehmann and Romano, 2005, Chapter 15.2). Hence, the null hypothesis can be simulated by permuting samples, and the permutation test can then be constructed by following the above construction with randomly sampled permutations. The resulting test can then be guaranteed to control the type I error non-asymptotically at the desired level $\alpha$ (Romano and Wolf, 2005a, Lemma 1, see also Kim and Schrab, 2023, Lemma 15). As explained in Sections 4.1, 5.1 and 6.1, the two-sample and independence problems can be framed as testing for exchangeability, but the goodness-of-fit one cannot.
@@
where $\varepsilon_{1},\dots,\varepsilon_{n}$ are i.i.d. Rademacher variables [^3] We focus on this setting, while the wild bootstrap can more generally be used with any distribution having mean zero and variance one (e.g., standard Gaussian)., i.e., each taking value $+1$ or $-1$ with probability $\nicefrac{{1}}{{2}}$. Under the null hypothesis, with some assumptions on the core function $h$, the asymptotic distribution of the wild bootstrapped statistic can be proven to be the same as the one of the original statistic (Chwialkowski et al., 2014, Theorem 1). So, the wild bootstrap can be used to simulate the null distribution asymptotically. Hence, following the test construction depicted above leads to a test which controls the type I error at level $\alpha$ asymptotically. For some frameworks testing for exchangeability and with particular choices of the core function, computing the wild bootstrapped statistic of Equation 3 can be equivalent to computing the statistic on permuted samples for some specific permutation (see Schrab et al., 2023, Appendix B for the MMD two-sample case, and Schrab et al., 2022b, Appendix F.1 for the HSIC independence case). Leveraging these results, we obtain non-asymptotic type I error control for these wild bootstrap hypothesis tests.
@@
If the framework considered is not equivalent to testing exchangeability, then the permutation method does not apply, and one should rely on the wild bootstrap approach and its asymptotic level control (e.g., goodness-of-fit testing in Section 6). When testing for exchangeability (e.g., two-sample and independence testing in Sections 4 and 5), a clear advantage of the permutation method is that it can be applied when using any statistic, while the wild bootstrap method only works for one-sample second-order $V$-statistics and its incomplete variants (i.e., not for the two-sample MMD V- and U-statistics (e.g., Schrab, 2025b, Equations 7 and 12) for the case $m\neq n$, and not for the full fourth-order HSIC V- and U-statistics (e.g., Schrab, 2025b, Equations 21 and 26)). So, when using a one-sample second-order statistic, both methods can be used, and for two-sample and independence testing, using a wild bootstrap is equivalent to using a subset of permutations (Schrab et al., 2023, Appendix B and Schrab et al., 2022b, Appendix F.1), so both methods benefit from non-asymptotic level guarantees in that setting. When using a complete one-sample second-order statistic (i.e., U-statistics or V-statistics), then either method can be used and perform similarly. When using an incomplete one-sample second-order statistic (see Schrab, 2025b, Section 3), using a wild bootstrap is highly preferable for computational reasons, as computing a permuted incomplete statistic likely results in having to evaluate the core function at new pairs of data points (not included in the original design but belonging to the permuted design).
@@ 1.3 Power of hypothesis testing
$\sup_{P_{1}\in{\mathcal{S}}_{1}}\mathbb{P}_{P_{1}}(\textrm{reject }{\mathcal{H}}_{0})\leavevmode\nobreak\ \geq\leavevmode\nobreak\ 1-\beta.$ (5)
@@
    Upper bound: For a given level $\alpha$ test, to determine the smallest separation rate $\rho$ for which uniform power holds (Equation 5).
@@
    Lower bound: To determine the largest separation rate $\rho$ such that no level $\alpha$ test can achieve uniform power as in Equation 5.
@@
$\int_{\mathbb{R}^{d}}\big{\|}\xi\big{\|}^{2s}_{2}\,\big{|}\widehat{f}(\xi)\big{|}^{2}\,\mathrm{d}\xi\leavevmode\nobreak\ \leq\leavevmode\nobreak\ (2\pi)^{d}$ (6)
where the Fourier transform is $\widehat{f}(\xi)\coloneqq\displaystyle\int_{\mathbb{R}^{d}}f(x)e^{-ix\!^{\top}\!\xi}\,\mathrm{d}x$ for all $\xi\in\mathbb{R}^{d}$. Intuitively, if the parameter $s$ were to be zero, then, by Plancherel’s theorem, Equation 6 would hold as long as $\|f\|_{L^{2}}\leq 1$ which would not impose any smoothness constraint. If the parameter $s$ is large, however, then the term $\|\xi\|_{2}^{2s}$ in Equation 6 ensures that the Fourier transform decays at a rapid rate, which imposes a smoothness requirement.
@@ 2 Adaptive methods: aggregation multiple testing and kernel pooling
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}\Big{(}T^{(k)}_{{\tilde{\alpha}}}\textrm{ rejects }{\mathcal{H}}_{0}\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ {\tilde{\alpha}}$ (7)
@@
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}\Big{(}T^{(k)}_{\alpha/|{\mathcal{K}}|}\textrm{ rejects }{\mathcal{H}}_{0}\textrm{ for some }k\in{\mathcal{K}}\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \alpha.$ (8)
@@
$\mathbb{P}_{P_{0}}\Big{(}T^{(k)}_{{\tilde{\alpha}}_{P_{0}}}\textrm{ rejects }{\mathcal{H}}_{0}\textrm{ for some }k\in{\mathcal{K}}\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \alpha\leavevmode\nobreak\ \textrm{ for all }\leavevmode\nobreak\ P_{0}\in{\mathcal{P}}_{0},$ (10)
@@
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}\Big{(}T^{(k)}_{{\tilde{\alpha}}_{P_{0}}}\textrm{ rejects }{\mathcal{H}}_{0}\textrm{ for some }k\in{\mathcal{K}}\Big{)}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \alpha.$ (11)
The probability $\mathbb{P}_{\pi(P)}$ in Equation 9 can be estimated with a Monte-Carlo procedure by permuting the samples (or via wild bootstrap), and the supremum in Equation 9 can be estimated using a bisection method. The aggregation level control of Equation 11 still holds non-asymptotically when using these estimated quantities (Schrab et al., 2023, Proposition 8). The aggregation procedure results in the most powerful multiple test which controls the type I error at level $\alpha$, this test always outperforms the multiple test with Bonferroni correction as $\tilde{\alpha}_{P}$ is always greater or equal to $\alpha$. Moreover, in Equation 9, it is possible to attribute some different weight $w_{k}$ (all summing to a quantity less or equal to one) to each kernel $k\in{\mathcal{K}}$ in the level correction $u$. See implementation details in Schrab et al. (2023, Algorithm 1 and Section 3.5). Working with the p-value view, the aggregation procedure can be linked to the method of Shah and Bühlmann (2018); detailed connections are left for future work.
@@
Another method to construct an adaptive test is to use an adaptive estimator in the first place. We call this kernel pooling (Schrab, 2025b, Section 4): given a collection of estimators $S_{k}$ for $k\in{\mathcal{K}}$, the pooled estimator is defined as $\textrm{pool}_{k\in{\mathcal{K}}}\leavevmode\nobreak\ S_{k}/\sigma_{k}$, where $\sigma_{k}$ is 1 for the unnormalised case, or is defined as in Schrab (2025b, Equation 73) for the normalised case. The pooling function ‘pool’ can be chosen to be the mean, the maximum, or fuse (Schrab, 2025b, Equations 74, 76 and 77). In practice, we recommend using the fuse pooling function, which is a soft maximum with a logsumpexp expression, as it empirically leads to a more powerful test. Fuse pooling is analysed in details in Biggs et al. (2023). The collection of kernels can be taken to be the same as the aforementioned one used for aggregation (e.g., Schrab, 2025b, Equation 81). To conclude, adaptive tests can be constructed either via aggregation or via kernel pooling.
@@ 3 Testing constraints: efficiency, privacy & robustness
(reject $\mathbb{P}\bigl{(}\textrm{reject }{\mathcal{H}}_{0}\textrm{ using}\mathbb{X}_{n}\bigr{)}\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ e^{\varepsilon}\,\mathbb{P}\bigl{(}\textrm{reject }{\mathcal{H}}_{0}\textrm{ using }\widetilde{\mathbb{X}}_{n}\bigr{)}+\delta,$ (12)
(fail to reject $\mathbb{P}\bigl{(}\textrm{fail to reject }{\mathcal{H}}_{0}\textrm{ using }\mathbb{X}_{n}\bigr{)}\leavevmode\nobreak$ $\leq\leavevmode\nobreak\ e^{\varepsilon}\,\mathbb{P}\bigl{(}\textrm{fail to reject }{\mathcal{H}}_{0}\textrm{ using }\widetilde{\mathbb{X}}_{n}\bigr{)}+\delta,$
@@
Another practical problem is the one of corrupted data, in real-world applications it is often the case that a portion of the data does not actually follow the distributions we would like to test (e.g., fake data, adversarially corrupted data, etc.). In some cases, these outliers are actually important and we want the tests to able to detect them. In other cases, these are just noise that we wish to be robust against in order to test the real problem. The aim of robust testing is then to test the null hypothesis when at most $r$ samples have been corrupted, possibly in an adversarial manner (this setting is more general than Huber’s contamination model). That is, the new ‘robust null hypothesis’ is that the condition ${\mathcal{H}}_{0}$ holds for at least $N-r$ of the $N$ samples, where the robustness parameter $r$ is specified by the user in advance depending on the testing setting and the amount of robustness required. See Schrab and Kim (2025, Section 2.1) for details on the robust testing framework. While we show that differentially private tests with adjusted level can be robust (Schrab and Kim, 2025, Algoritm 2), our main contribution is to propose a new procedure to robustify any permutation test (Schrab and Kim, 2025, Algoritm 1) by adding a factor of $2r\Delta$ to the rejection threshold (i.e., quantile obtained using permutations), where $\Delta$ is the global sensitivity of the test statistic (Kim and Schrab, 2023, Definition 2). Using these procedures, robust two-sample and independence tests, dcMMD and dcHSIC, can be constructed, which are robust up to $r$ corruption.

### v2

@@ Abstract
This paper provides a unifying view of optimal kernel hypothesis testing across the MMD two-sample, HSIC independence, and KSD goodness-of-fit frameworks. Minimax optimal separation rates in the kernel and $L^{2}$ metrics are presented, with two adaptive kernel selection methods (kernel pooling and aggregation), and under various testing constraints: computational efficiency, differential privacy, and robustness to data corruption. Intuition behind the derivation of the power results is provided in a unified way across the three frameworks, and open problems are highlighted.
@@
In Section˜1, we formalise the hypothesis testing framework, highlighting test level and power properties. In Section˜2, we present multiple testing via the aggregation method, which is strictly more powerful than using a Bonferroni correction. In Section˜3, we consider hypothesis testing under three constraints: computational efficiency, differential privacy and robustness to data corruption. In Sections˜4, 5 and 6, we introduce the two-sample, independence and goodness-of-fit testing frameworks, respectively, and derive power guarantees under kernel and $L^{2}$ uniform separation in the standard, efficient, private and robust settings, with aggregation and pooling kernel adaptation methods. In Section˜7, we highlight open problems which are left for future work. In Appendix˜A, we provide intuition behind the proofs of all the power results presented.
@@ 1 Hypothesis testing
We formally introduce the concept of hypothesis testing in Section˜1.1, we discuss test level control in Section˜1.2 and test power guarantees in Section˜1.3.
@@ 1.1 Definition of hypothesis testing
For two-sample testing, the space $\mathcal{P}$} consists of all pairs of distributions, and ${\mathcal{P}}_{0}$ of only the pairs with the same distribution for each component. For independence testing, the space $\mathcal{P}$} consists of joint distributions, and ${\mathcal{P}}_{0}$ of the ones which are equal to the product of their marginals. For goodness-of-fit testing, we have a reference model distribution $P_{\textrm{model}}$, the space $\mathcal{P}$} consists of all distributions while ${\mathcal{P}}_{0}=\{P_{\textrm{model}}\}$ is simply the model distribution. For more details on these testing frameworks, see Sections˜4, 5 and 6.
@@
A test is a function which takes as input the i.i.d. samples from $P\in{\mathcal{P}}$, and returns 0 if the null ${\mathcal{H}}_{0}$ is believed to hold, or 1 otherwise (i.e. ${\mathcal{H}}_{1}$ is believed to hold). A test is usually constructed by first computing a statistic, a real value computed from the data which is designed to capture evidence against the null when it exists. This statistic is then compared to a rejection region to decide whether to reject the null. Equivalently, the test can be defined via its p-value rather than via its quantile (see details below in Section˜1.2).
@@ 1.2 Level of hypothesis testing
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}(\textrm{reject }{\mathcal{H}}_{0})\penalty 10000\ \leq\penalty 10000\ \alpha.$ (2)
Typically, we would want this inequality to be tight: if the type I error is required to be less than $\alpha$, we would ideally want it to be as close as possible to $\alpha$ since a smaller type I error (e.g., conservative test) would result in a larger type II error (see Section˜1.3).
@@
A sample is said to be exchangeable if, for any permutation, the joint distribution over the permuted samples is equal to the joint distribution over the original samples. Some hypothesis testing problems can be framed as testing whether exchangeability holds (Lehmann and Romano, 2005, Chapter 15.2). Hence, the null hypothesis can be simulated by permuting samples, and the permutation test can then be constructed by following the above construction with randomly sampled permutations. The resulting test can then be guaranteed to control the type I error non-asymptotically at the desired level $\alpha$ (Romano and Wolf, 2005a, Lemma 1, see also Kim and Schrab, 2023, Lemma 15). As explained in Sections˜4.1, 5.1 and 6.1, the two-sample and independence problems can be framed as testing for exchangeability, but the goodness-of-fit one cannot.
@@
where $\varepsilon_{1},\dots,\varepsilon_{n}$ are i.i.d. Rademacher variables [^3] We focus on this setting, while the wild bootstrap can more generally be used with any distribution having mean zero and variance one (e.g., standard Gaussian)., i.e., each taking value +1+1 or 1-1 with probability $\nicefrac{{1}}{{2}}$. Under the null hypothesis, with some assumptions on the core function $h$, the asymptotic distribution of the wild bootstrapped statistic can be proven to be the same as the one of the original statistic (Chwialkowski et al., 2014, Theorem 1). So, the wild bootstrap can be used to simulate the null distribution asymptotically. Hence, following the test construction depicted above leads to a test which controls the type I error at level $\alpha$ asymptotically. For some frameworks testing for exchangeability and with particular choices of the core function, computing the wild bootstrapped statistic of Equation˜3 can be equivalent to computing the statistic on permuted samples for some specific permutation (see Schrab et al., 2023, Appendix B for the MMD two-sample case, and Schrab et al., 2022b, Appendix F.1 for the HSIC independence case). Leveraging these results, we obtain non-asymptotic type I error control for these wild bootstrap hypothesis tests.
@@
If the framework considered is not equivalent to testing exchangeability, then the permutation method does not apply, and one should rely on the wild bootstrap approach and its asymptotic level control (e.g., goodness-of-fit testing in Section˜6). When testing for exchangeability (e.g., two-sample and independence testing in Sections˜4 and 5), a clear advantage of the permutation method is that it can be applied when using any statistic, while the wild bootstrap method only works for one-sample second-order $V$-statistics and its incomplete variants (i.e., not for the two-sample MMD V- and U-statistics (e.g., Schrab, 2025b, Equations 7 and 12) for the case $m\neq n$, and not for the full fourth-order HSIC V- and U-statistics (e.g., Schrab, 2025b, Equations 21 and 26)). So, when using a one-sample second-order statistic, both methods can be used, and for two-sample and independence testing, using a wild bootstrap is equivalent to using a subset of permutations (Schrab et al., 2023, Appendix B and Schrab et al., 2022b, Appendix F.1), so both methods benefit from non-asymptotic level guarantees in that setting. When using a complete one-sample second-order statistic (i.e., U-statistics or V-statistics), then either method can be used and perform similarly. When using an incomplete one-sample second-order statistic (see Schrab, 2025b, Section 3), using a wild bootstrap is highly preferable for computational reasons, as computing a permuted incomplete statistic likely results in having to evaluate the core function at new pairs of data points (not included in the original design but belonging to the permuted design).
@@ 1.3 Power of hypothesis testing
$\sup_{P_{1}\in{\mathcal{S}}_{1}}\mathbb{P}_{P_{1}}(\textrm{reject }{\mathcal{H}}_{0})\penalty 10000\ \geq\penalty 10000\ 1-\beta.$ (5)
@@
    Upper bound: For a given level $\alpha$ test, to determine the smallest separation rate $\rho$ for which uniform power holds (Equation˜5).
@@
    Lower bound: To determine the largest separation rate $\rho$ such that no level $\alpha$ test can achieve uniform power as in Equation˜5.
@@
$\int_{\mathbb{R}^{d}}\big\|\xi\big\|^{2s}_{2}\,\big|\widehat{f}(\xi)\big|^{2}\,\mathrm{d}\xi\penalty 10000\ \leq\penalty 10000\ (2\pi)^{d}$ (6)
where the Fourier transform is $\widehat{f}(\xi)\coloneqq\displaystyle\int_{\mathbb{R}^{d}}f(x)e^{-ix\!^{\top}\!\xi}\,\mathrm{d}x$ for all $\xi\in\mathbb{R}^{d}$. Intuitively, if the parameter $s$ were to be zero, then, by Plancherel’s theorem, Equation˜6 would hold as long as $\|f\|_{L^{2}}\leq 1$ which would not impose any smoothness constraint. If the parameter $s$ is large, however, then the term $\|\xi\|_{2}^{2s}$ in Equation˜6 ensures that the Fourier transform decays at a rapid rate, which imposes a smoothness requirement.
@@ 2 Adaptive methods: aggregation multiple testing and kernel pooling
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}\Big(T^{(k)}_{{\tilde{\alpha}}}\textrm{ rejects }{\mathcal{H}}_{0}\Big)\penalty 10000\ \leq\penalty 10000\ {\tilde{\alpha}}$ (7)
@@
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}\Big(T^{(k)}_{\alpha/|{\mathcal{K}}|}\textrm{ rejects }{\mathcal{H}}_{0}\textrm{ for some }k\in{\mathcal{K}}\Big)\penalty 10000\ \leq\penalty 10000\ \alpha.$ (8)
@@
[…]
@@
$\sup_{P_{0}\in{\mathcal{P}}_{0}}\mathbb{P}_{P_{0}}\Big(T^{(k)}_{{\tilde{\alpha}}_{P_{0}}}\textrm{ rejects }{\mathcal{H}}_{0}\textrm{ for some }k\in{\mathcal{K}}\Big)\penalty 10000\ \leq\penalty 10000\ \alpha.$ (11)
The probability $\mathbb{P}_{\pi(P)}$ in Equation˜9 can be estimated with a Monte-Carlo procedure by permuting the samples (or via wild bootstrap), and the supremum in Equation˜9 can be estimated using a bisection method. The aggregation level control of Equation˜11 still holds non-asymptotically when using these estimated quantities (Schrab et al., 2023, Proposition 8). The aggregation procedure results in the most powerful multiple test which controls the type I error at level $\alpha$, this test always outperforms the multiple test with Bonferroni correction as $\tilde{\alpha}_{P}$ is always greater or equal to $\alpha$. Moreover, in Equation˜9, it is possible to attribute some different weight $w_{k}$ (all summing to a quantity less or equal to one) to each kernel $k\in{\mathcal{K}}$ in the level correction $u$. See implementation details in Schrab et al. (2023, Algorithm 1 and Section 3.5). Working with the p-value view, the aggregation procedure can be linked to the method of Shah and Bühlmann (2018); detailed connections are left for future work.
@@
Another method to construct an adaptive test is to use an adaptive estimator in the first place. We call this kernel pooling (Schrab, 2025b, Section 4): given a collection of estimators $S_{k}$ for $k\in{\mathcal{K}}$, the pooled estimator is defined as $\textrm{pool}_{k\in{\mathcal{K}}}\penalty 10000\ S_{k}/\sigma_{k}$, where $\sigma_{k}$ is 1 for the unnormalised case, or is defined as in Schrab (2025b, Equation 73) for the normalised case. The pooling function ‘pool’ can be chosen to be the mean, the maximum, or fuse (Schrab, 2025b, Equations 74, 76 and 77). In practice, we recommend using the fuse pooling function, which is a soft maximum with a logsumpexp expression, as it empirically leads to a more powerful test. Fuse pooling is analysed in details in Biggs et al. (2023). The collection of kernels can be taken to be the same as the aforementioned one used for aggregation (e.g., Schrab, 2025b, Equation 81). To conclude, adaptive tests can be constructed either via aggregation or via kernel pooling.
@@ 3 Testing constraints: efficiency, privacy & robustness
[…]
@@
Another practical problem is the one of corrupted data, in real-world applications it is often the case that a portion of the data does not actually follow the distributions we would like to test (e.g., fake data, adversarially corrupted data, etc.). In some cases, these outliers are actually important and we want the tests to able to detect them. In other cases, these are just noise that we wish to be robust against in order to test the real problem. The aim of robust testing is then to test the null hypothesis when at most $r$ samples have been corrupted, possibly in an adversarial manner (this setting is more general than Huber’s contamination model). That is, the new ‘robust null hypothesis’ is that the condition ${\mathcal{H}}_{0}$ holds for at least NrN-r of the $N$ samples, where the robustness parameter $r$ is specified by the user in advance depending on the testing setting and the amount of robustness required. See Schrab and Kim (2025, Section 2.1) for details on the robust testing framework. While we show that differentially private tests with adjusted level can be robust (Schrab and Kim, 2025, Algoritm 2), our main contribution is to propose a new procedure to robustify any permutation test (Schrab and Kim, 2025, Algoritm 1) by adding a factor of $2r\Delta$ to the rejection threshold (i.e., quantile obtained using permutations), where $\Delta$ is the global sensitivity of the test statistic (Kim and Schrab, 2023, Definition 2). Using these procedures, robust two-sample and independence tests, dcMMD and dcHSIC, can be constructed, which are robust up to $r$ corruption.